# A Survey on Optimization Studies of Group Centrality Metrics

### v1

@@ Abstract
Centrality metrics have become a popular concept in network science and optimization. Over the years, centrality has been used to assign importance and identify influential elements in various settings, including transportation, infrastructure, biological, and social networks, among others. That said, most of the literature has focused on nodal versions of centrality. Recently, group counterparts of centrality have started attracting scientific and practitioner interest. The identification of sets of nodes that are influential within a network is becoming increasingly more important. This is even more pronounced when these sets of nodes are required to induce a certain motif or structure. In this study, we review group centrality metrics from an operations research and optimization perspective for the first time. This is particularly interesting due to the rapid evolution and development of this area in the operations research community over the last decade. We first present a historical overview of how we have reached this point in the study of group centrality. We then discuss the different structures and motifs that appear prominently in the literature, alongside the techniques and methodologies that are popular. We finally present possible avenues and directions for future work, mainly in three areas: (i) probabilistic metrics to account for randomness along with stochastic optimization techniques; (ii) structures and relaxations that have not been yet studied; and (iii) new emerging applications that can take advantage of group centrality. Our survey offers a concise review of group centrality and its intersection with network analysis and optimization.
Keywords: Centrality metrics, group centrality, network analysis, network optimization, optimization methods
@@ 1 Introduction
Centrality is arguably the most popular concept studied in network science, graph theory, and the operations research (OR) communities. It primarily focuses on the significance, influence, and/or criticality of elements in a given network consisting of nodes and edges. Centrality metrics are inherently fundamental in comprehending complex systems from a network topology perspective. This is immediately visible from the large number of studies and applications that employ them.
@@
Applications of centrality are wide-reaching and highly interdisciplinary. As an example, from a supply chain perspective, centrality can indicate the importance of a particular supplier (or group of suppliers) to the success of the supply chain operations. It can help decision makers identify potential bottlenecks in the supply chain network (Borgatti and Li, 2009). Similarly, from an infrastructure network perspective, centrality might express the importance of connection points in ensuring a reliable and resilient network in the face of disruptive and cascading events (Stergiopoulos et al., 2015). In fact, choosing the appropriate centrality measure is a crucial step in every application. We refer the reader to see Chebotarev and Gubanov (2020) for an extensive discussion on how to choose the right centrality metric depending on the application.
@@
Node-based centrality metrics (nodal centrality) are indices that aim to quantify the importance of a single node in the grand scheme of a network. In the beginning, the notion of point centrality, a precursor to the more famous nowadays betweenness centrality, was introduced in the fundamental work by Bavelas Bavelas (1948). Motivated and inspired by this work, more and more researchers from a wide variety of fields started focusing on other forms of centrality, including but not limited to degree (Bródka et al., 2011), closeness (Okamoto et al., 2008), betweenness (Freeman, 1977, 1978; Riondato and Kornaropoulos, 2014) and its more modern adaptation to flow betweenness (Freeman et al., 1991), and eigenvector centrality (Bonacich, 1972; Maharani et al., 2014) have been widely studied, applied, and taught. In fact, several survey and review studies have focused on node-based centrality metrics in the literature, especially in the context of network resiliency (Wan et al., 2021), biological networks (Wang et al., 2022), wireless sensor networks (Jain and Reddy, 2013), and general complex networks (Saxena and Iyengar, 2020), among others.
@@
Group centrality metrics, on the other hand, focus not on individual nodes but rather on identifying clusters or groups of network elements that collectively maximize or minimize specific criteria (Everett and Borgatti, 1999). From an application standpoint, researchers have applied group centrality metrics to various domains, including social networks (Das et al., 2018), infrastructure networks (Michalak et al., 2013), biological networks (Vogiatzis and Camur, 2019), traffic networks (Ohara et al., 2017), human disease networks (Muhiuddin et al., 2023), and transportation networks (Vogiatzis and Pardalos, 2016). Overall, group centrality has become a vital tool in the analysis of networks, one that provides a unique perspective on network dynamics in various fields by analyzing the collective behavior of its interconnected elements.
@@
It is important to mention that there is growing interest in the OR community regarding primarily cohesive group (Frank, 1995) centrality metrics and their applications, especially in the last decade. Given a network, a cohesive group may represent the group of nodes which are more densely interconnected among themselves than with nodes outside the group. This, intentionally open-ended, definition allows for different interpretations in different contexts. However, to the best of our knowledge, no new/recent review study has been conducted on specifically group centrality metrics from an optimization perspective. Hence, our work intends to provide a concise survey of how group centrality metrics have been evolved over time and how they have been employed by OR researchers to develop new optimization models and methodologies.
@@ 2 Definitions and extensions
In this section, we provide details regarding the concepts used throughout the paper. Let us assume that we are given a simple, undirected, and connected (i.e., there exists a path between each part of nodes) network $G=(V,E)$. We refer the reader to (Freeman, 1977) for detailed discussion on centrality in disconnected graphs. As is common, $V=\{1,2,\cdots,n\}$ and $E\subseteq V\times V$ represent the set of nodes and edges, respectively. Given a node $v\in V$, let $N(v)$ be the set of nodes that are neighbors to $v$: mathematically, $N(v)=\left\{u:(u,v)\in E\right\}$.
Moreover, let $\mathcal{P}_{ut}$ be the set of all shortest (geodesic) paths connecting nodes $u$ and $t$ ($\mathcal{P}_{ut}\neq\emptyset$, due to connectedness). We use a similar notation for the number of shortest paths connecting $u$ to $t$ that pass through node $v\neq u,t$. Specifically, we let this quantity be $\sigma_{ut}(v)$. For simplicity, we let $\sigma_{ut}(\cdot)=|\mathcal{P}_{ut}|$.
@@
  • Degree centrality: The number of adjacent nodes to $v$ represented as $|N(v)|$. Often, we report the normalized version of that quantity as $\frac{|N(v)|}{|V|-1}$.
@@
  • Closeness centrality: The inverse of the sum of shortest path between $v$ and all other nodes and is denoted as $\sum_{u\in V}\frac{1}{d_{uv}}$.
Given a subset of nodes $S\subseteq V$, we define $G[S]$ as the induced subgraph where the edge set is $S\times S\cap E$. In this work, we use $G[S]$ for any group structure. Then, the centrality concepts introduced above can be extended to sets of nodes $S\subseteq V$ with a slight misuse of notation.
@@
We present Example 1 (based on Figure 1) showing the calculation for the three definitions above for a set of nodes that does not induce a connected subgraph.
@@
Finally, the group betweenness centrality is calculated as follows. First, we note how there are $132$ possible paths to consider (from node $1$ to nodes $3,5,7,\ldots,16$, and then from node $3$ to nodes $1,5,7,\ldots,16$, and so on). Then, we observe how paths originating from some nodes will always use at least one node in $S$ (nodes $1,5,7$). Others will sometimes use at least one node in $S$, whereas others will not. Specifically, look at node $3$: there are paths to $7$ nodes ($1,5,7,13,14,15,16$) that will always use at least one node in $S$ and another 4 paths to nodes $8,10,11,12$ that will not. The same calculation is true for nodes $8,10,11,12$. Finally, nodes $13,14,15,16$ will use at least one node in $S$ only for their paths towards $1,3,5,7,8,10,11,12$. In total, we have that the betweenness centrality of $S$ is calculated as $\frac{3\cdot 11+5\cdot 7+4\cdot 8}{132}=\frac{100}{132}$. ∎
@@
A set of nodes $C\subseteq V$ forms a clique if $C$ is inducing a complete subgraph (Pardalos and Xue, 1994). In other words, $C\subseteq V$ forms a clique if all edges $(i,j),i\in C,j\in C$ are present ($E[G[C]]=C\times C$). On the other hand, $C\subseteq V$ is said to form a star if there exists exactly one node $s\in C$ such that $\left(s,j\right)\in E$ for all other nodes $j\in C\setminus\left\{s\right\}$ and no other edge $\left(i,j\right)$ exists between any two nodes $i,j\in C\setminus\left\{s\right\}$. We observe how $s$ has a prominent role and is referred to as the center of the star, whereas all other nodes $i\in C\setminus\left\{s\right\}$ are referred to as the leaves of the star. We also note that both cliques and stars appear to have a relationship to independent sets, where any two nodes cannot be adjacent. A clique in graph $G(V,E)$ is an independent set in the complement graph $G^{\prime}(V,E^{\prime})$ where $\left(i,j\right)\in E^{\prime}$ if and only if $\left(i,j\right)\notin E$. The leaves in the star in $G(V,E)$ also induce an independent set in the same graph $G$.
@@
The diameter is the length/cost of the longest of all shortest paths in the network. In a clique, the diameter is equal to $1$ as all shortest paths are between immediately adjacent nodes. Relaxing the diameter restriction affects the group’s reachability. A group of nodes inducing a subgraph with diameter smaller than or equal to $k$ is referred to as a $k$-club (Mokken et al., 1979). The density of a graph is the fraction of edges that are present compared to all possible edges being present. The clique, being an ideal subgraph, has density equal to 1 as all possible edges are present. Relaxing this restriction allows some of the edges to be absent and affects the group’s familiarity. If fraction $\gamma\in\left[0,1\right]$ of all possible edges is present, then this group forms a $\gamma$-quasi-clique (Abello et al., 2002). As is easy to check $\gamma=1$ is the same as a clique, and $\gamma=0$ can be any group of nodes.
Next, we focus on degree (as defined earlier): clique members $i\in C$ always have a degree of $|C|-1$; relaxing this requirement to having a degree of at least $|C|-k$ leads to the definition of $k$-plex (Seidman and Foster, 1978; Balasundaram et al., 2011). We can also see another common clique relaxation for the members’ degrees: a $k$-core is defined as the induced subgraph where all nodes have minimum degree equal to $k$ (Seidman, 1983). Last, we look at connectivity (or the group’s robustness). To disconnect a clique, we would need to remove almost all of its members. Specifically, to produce a disconnected graph after removing members of a clique $C$, we would need to delete $|C|-1$ of them. On the other hand, relaxing this to disconnect after removal of $|C|-k$ nodes results in the definition of a $k$-block.
@@
This topic has been quite prominent in the optimization community, as we will discuss in the next Section. An example of how the previously defined three notions of group centrality change in the case of cliques is discussed next based on Figure 3 adapted from the works in (Vogiatzis et al., 2015) and (Rasti and Vogiatzis, 2022).
@@
Finally, the clique betweenness for them is equal to $\frac{19}{33}$ and $\frac{28}{55}$. For this calculation, we first need all shortest paths between any two nodes outside $S_{1}$ (and later outside $S_{2}$). There are 11 shortest paths from any node to all other nodes: that is, a shortest path to any node except for the three nodes in $S_{1}$. For nodes $1,5,6,7$ all of their shortest paths will use at least one node in $S$. On the other hand, the shortest paths originating from nodes $8,9,10,11,12,13,14,15$ do not necessarily use any node in $S$, unless the destination is one of nodes $1,5,6,7$. There is a total of $132$ shortest paths, and a total of $4\cdot 11+8\cdot 4=76$ use at least one node in $S$ as an intermediary. Based on this calculation, the betweenness centrality for $S_{1}$ is $\frac{76}{132}=\frac{19}{33}.$
Similar calculations lead to the betweenness of $S_{2}$. Specifically, node $12$ will always use at least one node in $S_{2}$; nodes $13,14,15$ will use at least one node in $S_{2}$ when the path terminates at any node in $\left\{1,2,3,4,5,6,7,12\right\}$; finally, nodes $1,2,3,4,5,6,7$ will use at least one node in $S_{2}$ for their shortest paths towards nodes $12,13,14,15$. Out of a total of $110$ shortest paths, $1\cdot 10+3\cdot 8+7\cdot 4=62$ will use at least one node in $S_{2}$ and hence its betweenness is $\frac{62}{110}=\frac{31}{55}$.
@@
Example 3.
Note a clique of size $3$ containing nodes $\left\{8,9,10\right\}$. It is not maximal, as it is a subset of the clique $S_{2}$ from Example 2. That said, it can (despite it not being maximal) have higher centrality than the maximal clique it belongs to. Specifically, in this example, its degree is $4$ (its open neighborhood consists of nodes $3,11,12,13$), which is larger than the degree of $S_{2}$. Additionally, it could have different closeness and betweenness values. ∎

### v3

# largesymbols"00 largesymbols"01

@@ Abstract
Centrality metrics have become a popular concept in network science and optimization. Over the years, centrality has been used to assign importance and identify influential elements in various settings, including transportation, infrastructure, biological, and social networks, among others. That said, most of the literature has focused on nodal versions of centrality. Recently, group counterparts of centrality have started attracting scientific and practitioner interest. The identification of sets of nodes that are influential within a network is becoming increasingly more important. This is even more pronounced when these sets of nodes are required to induce a certain motif or structure. In this study, we review group centrality metrics from an operations research and optimization perspective for the first time. This is particularly interesting due to the rapid evolution and development of this area in the operations research community over the last decade. We first present a historical overview of how we have reached this point in the study of group centrality. We then discuss the different structures and motifs that appear prominently in the literature, alongside the techniques and methodologies that are popular. We finally present possible avenues and directions for future work, mainly in three areas: (i) probabilistic metrics to account for randomness along with stochastic optimization techniques; (ii) structures and relaxations that have not been yet studied; and (iii) new emerging applications that can take advantage of group centrality. Our survey offers a concise review of group centrality and its intersection with network analysis and optimization. Keywords: Centrality metrics, group centrality, network analysis, network optimization, optimization methods, graph theory
@@ 1 Introduction
Centrality is one of the most extensively studied concepts in network science, graph theory, and operations research (OR) communities. It primarily focuses on the significance, influence, and/or criticality of elements in a given network consisting of nodes and edges. Centrality metrics are inherently fundamental in comprehending complex systems from a network topology perspective. This is immediately visible from the large number of studies and applications that employ them.
@@
Applications of centrality are wide-reaching and highly interdisciplinary. As an example, from a supply chain perspective, centrality can indicate the importance of a particular supplier (or group of suppliers) to the success of supply chain operations. It can help decision makers identify potential bottlenecks in the supply chain network [14]. Similarly, from an infrastructure network perspective, centrality might express the importance of connection points in ensuring a reliable and resilient network in the face of disruptive and cascading events [91]. In fact, choosing the appropriate centrality measure is a crucial step in every application. We refer the reader to see Chebotarev and Gubanov [21] for an extensive discussion on how to choose the right centrality metric depending on the application.
@@
Node-based centrality metrics (nodal centrality) are indices that aim to quantify the importance of a single node in the grand scheme of a network. In the beginning, the notion of point centrality, a precursor to the more famous nowadays betweenness centrality, was introduced in the fundamental work by Bavelas [8]. Motivated and inspired by this work, more and more researchers from a wide variety of fields started focusing on other forms of centrality, including but not limited to degree [17], closeness [74], betweenness [36, 37, 82] and its more modern adaptation to flow betweenness [38], and eigenvector centrality [12, 58] have been widely studied, applied, and taught. In fact, several survey and review studies have focused on node-based centrality metrics in the literature, especially in the context of network resiliency [105], biological networks [106], wireless sensor networks [48], and general complex networks [84], among others.
Group centrality metrics, on the other hand, focus not on individual nodes but rather on identifying clusters or groups of network elements that collectively maximize or minimize specific criteria [29]. They play a pivotal role as they help identify influential groups of nodes, which is more applicable in scenarios where collective influence or group structure is crucial. For example, in social networks, a group of interconnected individuals may have a greater combined influence on information spread than the same individuals considered separately. In infrastructure networks, such as transportation or utility networks, group centrality can identify critical clusters of components whose simultaneous failure could lead to significant disruptions. This perspective is not captured by node centrality, which only assesses the importance of individual components.
@@
From an application standpoint, researchers have applied group centrality metrics to various domains, including social networks [24], infrastructure networks [64], biological networks [99], traffic networks [73], human disease networks [67], and transportation networks [101]. Overall, group centrality has become a vital tool in the analysis of networks, one that provides a unique perspective on network dynamics in various fields by analyzing the collective behavior of its interconnected elements.
@@
It is important to mention that there is growing interest in the OR community regarding primarily cohesive group [35] centrality metrics and their applications, especially in the last decade. Given a network, a cohesive group may represent the group of nodes which are more densely interconnected among themselves than with nodes outside the group. This, intentionally open-ended, definition allows for different interpretations in different contexts.
We note that group centrality is examined as a sub category under centrality measures in different survey papers (see Saxena and Iyengar [84] and Wan et al. [105].) However, to the best of our knowledge, no review study has been conducted on specifically group centrality metrics from an optimization and OR perspective. Hence, our work intends to provide a concise survey of how group centrality metrics have been evolved over time and how they have been employed by OR researchers to develop new optimization models and methodologies.
@@ 2 Definitions and extensions
In this section, we provide details regarding the concepts used throughout the paper. Let us assume that we are given a simple, undirected, and connected (i.e., there exists a path between each part of nodes) network $G=\left(V,E\right)$. We refer the reader to Freeman [36] for detailed discussion on centrality in disconnected graphs. As is common, $V=\{1,2,\cdots,n\}$ and $E\subseteq V\times V$ represent the set of nodes and edges, respectively. Given a node $v\in V$, let $N(v)$ be the set of nodes that are neighbors to $v$: mathematically, $N(v)=\left\{u:(u,v)\in E\right\}$.
Moreover, let $\mathcal{P}_{ut}$ be the set of all shortest (geodesic) paths connecting nodes $u$ and $t$ ($\mathcal{P}_{ut}\neq\emptyset$, due to connectedness). We use a similar notation for the number of shortest paths connecting $u$ to $t$ that pass through node $v\neq u,t$. Specifically, we let this quantity be $\sigma_{ut}(v)$. For simplicity, we let $\sigma_{ut}(\cdot)=\left|\mathcal{P}_{ut}\right|$.
@@
  • Degree centrality: The number of adjacent nodes to $v$ represented as $\left|N\left(v\right)\right|$. Often, we report the normalized version of that quantity as $\frac{\left|N\left(v\right)\right|}{\left|V\right|-1}$.
@@
  • Closeness centrality: The inverse of the sum of shortest path between $v$ and all other nodes and is denoted as $\sum_{u\in V}\frac{1}{d_{uv}}$. We note here that some works calculate closeness centrality as $\frac{1}{\sum_{u\in V}d_{uv}}$.
The three above centrality metrics [39] have been traditionally used as the basis of development for other variants of centrality. Specifically, betweenness and closeness are both based on shortest paths in a network. However, as noted in [90], using shortest paths can sometimes be limiting. Calculating then betweenness and closeness using all paths (or all paths bounded by some distance threshold) leads to the development of information or current-flow betweenness and closeness centrality metrics.
Given a subset of nodes $S\subseteq V$, we define $G\left[S\right]$ as the induced subgraph where the edge set is $S\times S\cap E$. In this work, we use $G\left[S\right]$ for any group structure. Then, the centrality concepts introduced above can be extended to sets of nodes $S\subseteq V$ with a slight misuse of notation.
@@
In some instances (like in closeness centrality), it may make sense to define a bottleneck version (or, equivalently, maximum distance closeness centrality). In that version, the biggest difference is that we are focused on the worst-case possible distance from any node outside $S$ to some node in $S$. In mathematical terms, the maximum distance closeness centrality is defined as $1/\max_{u\in V}d_{uS}$, where $d_{uS}=\min_{v\in S}d_{uv}$. We present Example 1 (based on Figure 1) showing the calculation for the three definitions above for a set of nodes that does not induce a connected subgraph.
@@
Finally, the group betweenness centrality is calculated as follows. First, we will consider nodes in lexicographic order, for simplicity. That is, starting from node $1$, we observe that there is exactly one shortest path towards nodes $3$, $5$, $7$, $8$, $10$, $11$, $13$, $14$, $15$, and $16$ and it always uses at least one node in $S$; furthermore, it is only node $12$ that has three shortest paths originating from $1$, however all three again use at least one node in $S$. This brings us to a total of 11. Continuing to node $3$, we have 2 shortest paths towards $5$ with both using a node in $S$ and 1 shortest path towards $7$ that also uses a node in $S$. Similarly, we have only one shortest path towards nodes $8,10,11$ which are not passing through a node in $S$, as well as one shortest path towards nodes $13,14,15,16$, all of which are passing through a node in $S$. We left the shortest paths from node $3$ to node $12$ for last: there are two shortest paths, only one of which is passing through a node in $S$ (namely, node $6$). Summing up all the fractions, we have a total of $7.5$.
Moving on, we can calculate that same count originating from nodes $5$ (9, when excluding shortest paths towards $1$ and $3$) and $7$ (8, excluding shortest paths towards $1,3,5$). Nodes $8,10,11,12$ have summations equal to $4$. This brings our total to $11+6.5+9+8+4+4+4+4=50.5$.
All of the calculations for betweenness thus far are not normalized. A normalized value can be computed by dividing the obtained score by $\frac{(|V|-|S|)\cdot(|V|-|S|-1)}{2}$. In our previous calculation, where $|V|=16$ and $|S|=4$, the normalized calculation would lead to $50.5/66=0.765$
@@
A set of nodes $C\subseteq V$ forms a clique if $C$ is inducing a complete subgraph [75]. In other words, $C\subseteq V$ forms a clique if all edges $(i,j),i\in C,j\in C$ are present ($E[G[C]]=C\times C$). On the other hand, $C\subseteq V$ is said to form a star if there exists exactly one node $s\in C$ such that $\left(s,j\right)\in E$ for all other nodes $j\in C\setminus\left\{s\right\}$ and no other edge $\left(i,j\right)$ exists between any two nodes $i,j\in C\setminus\left\{s\right\}$. We observe how $s$ has a prominent role and is referred to as the center of the star, whereas all other nodes $i\in C\setminus\left\{s\right\}$ are referred to as the leaves of the star. We also note that both cliques and stars appear to have a relationship to independent sets, where any two nodes cannot be adjacent. A clique in graph $G\left(V,E\right)$ is an independent set in the complement graph $G^{\prime}\left(V,E^{\prime}\right)$ where $\left(i,j\right)\in E^{\prime}$ if and only if $\left(i,j\right)\notin E$. The leaves in the star in $G\left(V,E\right)$ also induce an independent set within $G$.
@@
The diameter is the length/cost of the longest of all shortest paths in the network. In a clique, the diameter is equal to $1$ as all shortest paths are between immediately adjacent nodes. Relaxing the diameter restriction affects the group’s reachability. A group of nodes inducing a subgraph with diameter smaller than or equal to $k$ is referred to as a $k$-club [66]. The density of a graph is the fraction of edges that are present compared to all possible edges being present. The clique, being an ideal subgraph, has density equal to 1 as all possible edges are present. Relaxing this restriction allows some of the edges to be absent and affects the group’s familiarity. If fraction $\gamma\in\left[0,1\right]$ of all possible edges is present, then this group forms a $\gamma$-quasi-clique [1]. As is easy to check $\gamma=1$ is the same as a clique, and $\gamma=0$ can be any group of nodes.
Next, we focus on degree (as defined earlier): clique members $i\in C$ always have a degree of $|C|-1$; relaxing this requirement to having a degree of at least $|C|-k$ leads to the definition of $k$-plex [86, 7]. We can also see another common clique relaxation for the members’ degrees: a $k$-core is defined as the induced subgraph where all nodes have minimum degree equal to $k$ [85]. Last, we look at connectivity (or the group’s robustness). To disconnect a clique, we would need to remove almost all of its members. Specifically, to produce a disconnected graph after removing members of a clique $C$, we would need to delete $|C|-1$ of them. On the other hand, relaxing this to disconnect after removal of $|C|-k$ nodes results in the definition of a $k$-block.
@@
This topic has been quite prominent in the optimization community, as we will discuss in the next Section. An example of how the previously defined three notions of group centrality change in the case of cliques is discussed next based on Figure 3 adapted from the works in Vogiatzis et al. [102] and Rasti and Vogiatzis [81].
@@
Finally, the clique betweenness for them is equal to $38$ and $31$, respectively. For this calculation, we first need all shortest paths between any two nodes outside $S_{1}$ (and later outside $S_{2}$). By construction of the graph, there are 11 shortest paths from any node to all other nodes: that is, a shortest path to any node except for the three nodes in $S_{1}$. For nodes $1,5,6,7$ all of their shortest paths will use at least one node in $S$. On the other hand, the shortest paths originating from nodes $8,9,10,11,12,13,14,15$ do not use any node in $S$, unless the destination is one of nodes $1,5,6,7$, which have already been accounted for. Using these, we have a total betweenness score of $11+10+9+8=38$.
Similar calculations lead to the betweenness of $S_{2}$. Specifically, node $12$ will always use at least one node in $S_{2}$; nodes $13,14,15$ will use at least one node in $S_{2}$ when the path terminates at any node in $\left\{1,2,3,4,5,6,7,12\right\}$; finally, nodes $1,2,3,4,5,6,7$ will use at least one node in $S_{2}$ for their shortest paths only towards nodes $12,13,14,15$. Using these, we have a final score of $4+4+4+4+4+4+4+3+31$.
Once again, a normalized value can be calculated by dividing by the total number of pairs considered, which would be $66$ and $55$ for $S_{1},S_{2}$, respectively. Hence, the normalized values would be equal to $38/66=0.576$ and $31/55=0.564$
@@
Example 3
Note a clique of size $3$ containing nodes $\left\{8,9,10\right\}$. It is not maximal, as it is a subset of the clique $S_{2}$ from Example 2. That said, it can (despite it not being maximal) have higher centrality than the maximal clique it belongs to. Specifically, in this example, its degree is $4$ (its open neighborhood consists of nodes $3,11,12,13$), which is larger than the degree of $S_{2}$. Additionally, it could have different closeness and betweenness values.